OpenStudy (anonymous):

The length of a rectangle is four times its width. If the perimeter of the rectangle is 80cm , find its area.

OpenStudy (jdoe0001):

do you know what the Area and Perimeter formulas for a rectangle are?

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

well, then all you need to know is that the Length is 4 times the Width, that is, if Width is "W", Length is "4W", thus the perimeter will be 2(4W+W) = 80 from there you can solve for W to see what "W" is, once you have W, 4W gives you L when you have W and L, well, just use the Area formula :)

OpenStudy (anonymous):

let width=x length=4x perimeter=2(l+w)=2(4x+x)=10x 10x=80 x=8 width=8cm length=4*8=32cm area=32*8=256cm^2